OpenStudy (anonymous):

A pupil's marks were wrongly entered as 83 instead of 63.. Due to that the average marks for the class got increased by half. The number of pupils in the class are: a) 10 b) 20 c) 40 d) 73 I only require the solution and not the answer... Please help me with this...

OpenStudy (kropot72):

Let n = number of pupils Let x = incorrect average marks Incorrect total marks = x * n Corrected total marks = (x * n) - (83 - 63) = nx - 20 .................(1) Also the correct total marks = (x - 0.5)n = nx - 0.5n .................(2) Therefore nx - 20 = nx - 0.5n..................(3) Now just solve (3) to find n

OpenStudy (anonymous):

-20 = -0.5n n = 20/0.5 = 40...

OpenStudy (anonymous):

Okay, thanks buddy....

OpenStudy (kropot72):

You're welcome :)